# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v1

@@ 2.2 Persistent Homology on brain connectome
Persistent homology (PH) is a widely used effective algebraic topological tool for the analysis of the brain connectome, enabling the capture of more interpretable high-order topological features. Rather than attempting to determine a single optimal threshold, researchers (Hyekyoung et al., 2011) suggest examining the topological changes in the brain network as the threshold is increased continuously based on persistent homology. Researchers (Sizemore et al., 2018) also highlight the importance of cliques and cavities in the human connectome and locate topological cavities of different dimensions, around which information may flow in either diverging or converging patterns. To address the difficult issue of functional connectivity network (FCN) comparison across different spatiotemporal resolutions, researchers (Cassidy et al., 2018) develop a novel network comparison framework based on persistent homology to observe the change of 0-dimensional homology group. Besides, researchers (Bian et al., 2024) propose an adversarially trained persistent homology-based graph convolutional network (ATPGCN) to capture disease-specific brain connectome patterns for brain disorder diagnosis. Researchers (Bhattacharya et al., 2025) use the persistent topological features of connected components (H0) and loops (H1), which are quantified by two methods: Vietoris-Rips filtration and graph filtration, for diagnosing brain disorders. However, these studies ignore the importance of 2-dimensional topological profiles (i.e., voids), and do not links HOIs in the brain to these higher-order organisations. Recently, researchers (Santoro et al., 2024) have adopted certain methodologies grounded in the synchronisation phenomenon to construct more accurate simplicial complexes for modeling HOIs and use PH to capture 1-dimensional cycles related to triplet interactions. However, this can easily generate a large number of false positive high-order interactions and overlook higher-order quadruplet interactions with adequate temporal resolution. Besides, the influence of signed HOIs was overlooked in these studies, which could offer a broader diagnostic view of neurological conditions.
@@ 3.1 Signed Higher-Order Interaction Representation Strategy
Most traditional network neuroscience approaches construct functional connectivity through pairwise Pearson correlations of fMRI time series (Fornito et al., 2016; Lindquist, 2008). Recent edge-centric frameworks account for interactions of edges by computing instantaneous co-fluctuations through element-wise products of z-scored signals $\mathbf{x}_{i}\odot\mathbf{x}_{j}$ (Faskowitz et al., 2020). Reference (Santoro et al., 2023, 2024) extended this approach to characterize $k$-order interactions when $k+1$ nodes simultaneously deviate from their temporal mean activation at time $t$. Although these techniques based on Pearson correlations have provided insights into the brain’s dynamic higher-order interactions, the methods have generally been limited by the lack of adequate temporal resolution (Shine et al., 2015). Now we give a new way to calculate co-fluctuations between several ROIs.
@@ 3.1.3 Construct two signed weighted simplicial complexes
We must note that in a weighted simplicial complex $\mathcal{K}^{t}$, there are two kinds of signs of a simplex, the concordant signs of positive and the discordant signs of negative. We focus only on the concordant signs of positive but not on the discordant signs of negative because they may represent a confusing redundancy of information across multiple brain regions (Santoro et al., 2023). We further categorized the concordant interactions into positively-synergistic interactions and negatively-synergistic interactions. Positively synergistic interactions reflect multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment, which can offer valuable information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis.As shown in Fig. 2, it displays six different signed quadruplet higher-order interactions. For each temporal instance $t$, we further extract positively- and negatively-synergistic interactions from the weighted simplicial complexes to construct two signed weighted simplicial complexes $\mathcal{K}^{t}_{positive}$ and $\mathcal{K}^{t}_{negative}$ ,respectively.
@@ 3.3 Multi-Channel Brain Network Transformer
3.3.1 Signed High-order Features Decoupling Mechanism
High-order features that are positive are related to patterns of simultaneous activation in multiple brain regions at a given moment, while high-order features that are negative are related to patterns of collective inhibition in these regions at the current moment. These features can offer valuable information about complex coordination and communication within the brain. We further explore the influence of signed higher-order features using high-order features decoupling mechanism. Specifically, for each type of high-order feature’s two signed weighted adjacency matrix $A^{q}_{i},A^{p}_{i}\in\mathbb{R}^{N\times N}$ ($i=1,2$), we extract its upper triangular part and flatten it to construct a feature vector $h^{q}_{i},h^{p}_{i}\in\mathbb{R}^{N(N-1)/2}$. We generate two new high-order topological features by adaptively learning weights to aggregate positive and negative information:
@@ 3.3.2 Multi-Head Self-Attention Module of Multiple Channels
The feature matrixs extracted above actually represents the low-order and high-order topological information associated with each brain region. Simply concatenating the low-order and high-order information of each brain region as the input to the model would overlook the heterogeneity of the related information. Therefore, inspired by Brain network Transformer (Kan et al., 2022b), we leverage a $L$-layer non-linear mapping module, namely Multi-Head Self-Attention (MHSA), to generate more expressive node features using the low-order and high-order topological features of multiple channels ${Z}^{L}_{i}=\text{MHSA}({A_{i}})\in\mathbb{R}^{N\times N},(i=1,2,3)$. For each layer $l$, the output of each channel ${Z_{{i}}}^{l}$ is obtained by
@@ 4.1.2 Methods for comparison
The selected baselines correspond to five categories. The first category is traditional machine learning models by the scikit-learn library (Pedregosa et al., 2011) using edge features, including Multilayer Perceptron (MLP) , Support Vector Machine Classifier (SVM), Logistic Regression, and Random Forest. The second category consists of GNNs-based models that learn lower-order topological features from a two-way network. These models include GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024). The third category consists of Transformer-based models that utilize the Transformer architecture to perceive the entire brain picture, including Graph Transformer (Ying et al., 2021), Brain Network Transformer (Kan et al., 2022b), Tsen (Hu et al., 2023), Long-range Brain Transformer (Yu et al., 2024). The fourth category is high-order methods based on hypergraph, including HGCN (Hao et al., 2023), HGAT (Wang et al., 2022). The fifth category is persistent homology (PH)-based models, including Brain-HORS (Santoro et al., 2024), PH-MCI (Bhattacharya et al., 2025), ATPGCN (Bian et al., 2024).
@@ 4.1.3 Implementation Details
    For GNNs-based models and Transformer-based models, the correlation matrix is calculated using Pearson correlation for all time series. On this basis, the node features in the graph are the columns of the corresponding Pearson correlation matrix. The adjacency matrix in the graph is obtained by thresholding (0.5) the correlation matrix. Generally, two graph convolutional layers and an average pooling layer are used for classification.
@@
    For hypergraph-based high-order methods, specifically, for HGCN, a graph is first constructed through the correlation matrix, where the correlation of each edge of the graph is greater than 0.7. Each node constructs 2-hyperedges, 3-hyperedges, and 4-hyperedges by selecting its k-hop neighbors in the graph. For HGAT, the feature of the central node corresponding to each hyperedge is used as the hyperedge feature. When aggregating information, the attention coefficient between the hyperedge and the node is calculated for classification. Note that more HGNNs-based methods applying to brain data have been proposed recently. However, the code of these models is notpublicly available, and they require significant engineering which is hard to reproduce fairly based solely on the papers.
@@
    For the persistent homology (PH)-based models, including PH-MCI, ATPGCN. We use the persistent topological features of connected components (H0) and loops(H1) extracting by Ripserer tool. For Brain-HORS, following the method, we extract the matrix of violating triangles, the matrix of one-dimensional loops and the matrix of lower-order edge features, then flatten them and feed them into the SVM for classification.
@@
HGNNs-based models HGCN 58.9$\pm$7.6 61.2$\pm$7.0 60.1$\pm$7.7 59.5$\pm$7.3
@@
HGNNs-based models HGCN 55.0$\pm$20.3 58.7$\pm$23.1 55.0$\pm$20.3 52.4$\pm$20.2
@@
HGNNs-based models HGCN 64.2$\pm$10.0 65.2$\pm$10.3 64.3$\pm$10.3 63.4$\pm$10.5
@@
HGNNs-based models HGCN 61.0$\pm$3.0 55.2$\pm$5.3 62.3$\pm$4.6 60.4$\pm$4.3
@@ 4.2 Model Comparison
Table,2, 3, 4, 5 demonstrate the comparison results on ADNI, TaoWu, PPMI, and ABIDE datasets. The mean and standard deviation of ten-fold cross-validation were listed, with the best results highlighted in bold and the second-best results in underlining. Compared with traditional ML methods, general-purposed GNNs-based models for brain network do not show a significant advantage and attain similar performance in most cases. This finding is consistent with the results reported in existing papers. This is because, under the same input conditions, neural network models typically have a larger number of parameters than conventional machine learning methods when utilizing edge features for better interpretability, making them more susceptible to overfitting. Transformer-based models predominantly outperform GNN-based models across all four datasets in most scenarios, highlighting their superior capability to capture the global structure of brain connectomes. HGNNs-based models generally do not demonstrate superiority over other models, with notable performance only observed on the PPMI dataset. This phenomenon can be attributed to two main reasons: first, such models introduce excessive parameters, posing a risk of overfitting; second, their K-hop neighbor-based modeling approach fails to effectively capture the co-fluctuation characteristics among multiple brain regions, thereby inadequately representing high-order interactions. In PH-based models, such as ATPGCN and Brain-HORs, which introduce high-order loops or connected component features extracted via persistent homology for brain disease diagnosis, demonstrate significant superiority compared to both graph neural network (GNN)-based models and traditional machine learning models that rely solely on edge features. This highlights the necessity of incorporating high-order topological features in brain disease diagnosis methods and provides a theoretical foundation for exploring brain diseases from higher-dimensional and more comprehensive perspectives.
@@
Our HOI-brain outperforms all 20 baselines on all datasets. In particular, it can be found that the proposed HOI-brain outperforms the traditional machine learning models on all datasets, with average accuracy improved by 12.5 (%), average precision improved by 12.9(%), average recall improved by 14.3(%) and F1-score improved by 12.7(%), respectively. And the proposed HOI-brain outperforms the GNNs-based models on all datasets, with average accuracy improved by 19.6(%), average precision improved by 22.9(%), average recall improved by 19.5(%), and F1-score improved by 19.5(%), respectively. This indicates that our model, by incorporating high-order topological features, is more beneficial for brain disease diagnosis compared to using only low-order edge features. Furthermore, the proposed HOI-Brain outperforms Transformer-based models on all datasets, with average accuracy improved by 13.0(%), average precision improved by 13.4(%), average recall improved by 10.3(%), and F1-score improved by 10.9(%), respectively. This highlights that our model, through the multi-channel brain network transformer, can effectively and holistically integrate complementary information from heterogeneous high-order and low-order topological features while preserving channel-specific properties. Next, when compared to HGNNs-based models, HOI-Brain achieves remarkable improvements with average accuracy improved by 21.8(%), precision improved by 27.0(%), recall improved by 19.4(%), and F1-score improved by 20.5(%). These results clearly demonstrate that our proposed HOI-Brain framework significantly advances the modeling of higher-order interactions through the Multiplication of Temporal Derivatives and the extraction of topological features using persistent homology in brain networks. When specifically compared to Persistent Homology (PH) based models, HOI-Brain achieves significant improvements with average accuracy increased by 17.4(%), precision enhanced by 21.4(%), recall boosted by 14.5(%), and F1-score elevated by 16.2(%). This is partly attributed to our method linking signed HOIs in the brain to higher-order organizations, and partly due to our exploration of more complex signed quadruplet structures and voids (2D holes) formed by triangular faces and quadruplet.
@@ 4.3.1 Ablation study
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of different components on overall performance. Our first aim was to evaluate the effectiveness of more complex signed quadriplet structures and voids (2D holes) formed by triangular faces. Specifically, following the method in (Santoro et al., 2024), we extract lower-order edge features, violating triangles, one-dimensional loops. In addition, we extract unsigned good quadruplets and unsigned two-dimensional voids based our methods. We implement it within different combinations of these features on all datasets. The experimental results presented in Table 6, 7, 8, 9, indicate that integrating both higher-order and lower-order interactions proves to be more effective for brain disorder diagnosis compared to using lower-order interactions alone. Furthermore, quadruplet interactions captured by HOI-Brain have been demonstrated to be more effective in improving diagnosis than triplet interactions. This may be due to the fact that a significant proportion of triplet interactions can be decomposed into a linear combination of pairwise interactions, which is insufficient to capture the true high-order interactions. Additionally, distinguishing between positively and negatively synergistic higher-order interactions to generate signed higher-order structures, as opposed to unsigned ones, can further improve the effectiveness of brain disease diagnosis, which may due to the fact that it can offer detailed insights into the communication within the brain.
@@
Our second aim was to evaluate the effectiveness of different components in multi-channel brain network transformer. The different variants are: not adding signed high-order features decoupling mechanism (wo-signed), not adding orthonormal clustering readout of multiple channels, and not adding attention-guided feature fusion mechanism (wo-fusion), and not adding orthonormal clustering readout of multiple channels(wo-cluster). From Fig.4, we can see that adding signed high-order features decoupling mechanism and attention-guided feature fusion mechanism are always helpful as HOI-brain wins all the cases over the variant (wo-signed) or (wo-fusion). This result is consistent with our assumption that, by introducing the attention mechanism, our model can adaptively fuse positive and negative information as well as information from different channels, thereby emphasizing discriminative features while suppressing redundant information. Incorporating orthonormal clustering readout across multiple channels (wo-cluster) is highly beneficial, as this approach effectively captures both low-level and high-level feature patterns, which often reside within distinct functional modular structures.

### v2

@@ 2.2 Persistent Homology on brain connectome
Persistent homology (PH) is a widely utilised, effective algebraic topological tool for the analysis of the brain connectome, enabling the capture of more interpretable, higher-order topological features. In lieu of endeavouring to ascertain a solitary optimum threshold, researchers (Hyekyoung et al., 2011) propose the examination of the topological changes in the brain network as the threshold is increased continuously, based on persistent homology. Researchers (Sizemore et al., 2018) also highlight the importance of cliques and cavities in the human connectome, and locate topological cavities of different dimensions, around which information may flow in either diverging or converging patterns. In order to address the challenging issue of comparing functional connectivity networks (FCNs) across different spatiotemporal resolutions, researchers (Cassidy et al., 2018) have developed a novel network comparison framework based on persistent homology. The purpose of this framework is to observe the change of 0-dimensional homology groups. Furthermore, researchers (Bian et al., 2024) propose an adversarially trained persistent homology-based graph convolutional network (ATPGCN) to capture disease-specific brain connectome patterns for brain disorder diagnosis. Researchers (Bhattacharya et al., 2025) utilise the persistent topological features of connected components $H_{0}$ and loops $H_{1}$, which are quantified by two methods: The utilisation of Vietoris-Rips filtration and graph filtration is a pivotal diagnostic tool in the identification of brain disorders. Nevertheless, the aforementioned studies have overlooked the significance of two-dimensional topological profiles (i.e., voids) and have failed to establish a correlation between HOIs in the brain and these higher-order organisations. Recently, researchers (Santoro et al., 2024) have adopted certain methodologies grounded in the synchronisation phenomenon to construct more accurate simplicial complexes for modeling HOIs and use PH to capture 1-dimensional cycles related to triplet interactions. However, this can easily generate a large number of false positive higher-order interactions and overlook higher-order quadruplet interactions with adequate temporal resolution. Besides, the influence of signed HOIs was overlooked in these studies, which could offer a broader diagnostic view of neurological conditions.
@@ 3.1 Signed Higher-Order Interaction Representation Strategy
Most traditional network neuroscience approaches construct functional connectivity through pairwise Pearson correlations of fMRI time series (Fornito et al., 2016; Lindquist, 2008). Recent edge-centric frameworks account for interactions of edges by computing instantaneous co-fluctuations through element-wise products of z-scored signals $\mathbf{x}_{i}\odot\mathbf{x}_{j}$ (Faskowitz et al., 2020). Reference (Santoro et al., 2023, 2024) extended this approach to characterize $k$-order interactions when $k+1$ nodes simultaneously deviate from their temporal mean activation at time $t$. Although these techniques based on Pearson correlation have provided insights into the brain’s dynamic higher-order interactions, the methods have generally been limited by the lack of adequate temporal resolution (Shine et al., 2015). Now we give a new way to calculate co-fluctuations between several ROIs.
@@ 3.1.3 Construct two signed weighted simplicial complexes
It is important to acknowledge that in a weighted simplicial complex, denoted by $\mathcal{K}^{t}$, there exist two categories of signs for a simplex: concordant signs of positive and discordant signs of negative. The present study focuses exclusively on concordant signs of positive, eschewing the examination of discordant signs of negative. This is due to the potential for such signs to represent a confusing redundancy of information across multiple brain regions (Santoro et al., 2023). The concordant interactions were subsequently categorised into two distinct groups: those that were positively synergistic and those that were negatively synergistic. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer valuable insights into the complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. As demonstrated in Fig. 2, it exhibits six distinct signed quadruplet higher-order interactions. For each temporal instance $t$, the weighted simplicial complexes are further extracted to construct two signed weighted simplicial complexes $\mathcal{K}^{t}_{positive}$ and $\mathcal{K}^{t}_{negative}$, namely the positive and negative complexes, respectively.
@@ 3.3 Multi-Channel Brain Network Transformer
3.3.1 Signed higher-order Features Decoupling Mechanism
Higher-order features that are positive are related to patterns of simultaneous activation in multiple brain regions at a given moment, while higher-order features that are negative are related to patterns of collective inhibition in these regions at the current moment. These features can offer valuable information about complex coordination and communication within the brain. We further explore the influence of signed higher-order features using higher-order features decoupling mechanism. Specifically, for each type of higher-order feature’s two signed weighted adjacency matrix $A^{q}_{i},A^{p}_{i}\in\mathbb{R}^{N\times N}$ ($i=1,2$), we extract its upper triangular part and flatten it to construct a feature vector $h^{q}_{i},h^{p}_{i}\in\mathbb{R}^{N(N-1)/2}$. We generate two new higher-order topological features by adaptively learning weights to aggregate positive and negative information:
@@ 3.3.2 Multi-Head Self-Attention Module of Multiple Channels
The feature matrices extracted above actually represents the low-order and higher-order topological information associated with each brain region. Simply concatenating the low-order and higher-order information of each brain region as the input to the model would overlook the heterogeneity of the related information. Therefore, inspired by Brain network Transformer (Kan et al., 2022b), we leverage a $L$-layer non-linear mapping module, namely Multi-Head Self-Attention (MHSA), to generate more expressive node features using the low-order and higher-order topological features of multiple channels ${Z}^{L}_{i}=\text{MHSA}({A_{i}})\in\mathbb{R}^{N\times N},(i=1,2,3)$. For each layer $l$, the output of each channel ${Z_{{i}}}^{l}$ is obtained by
@@ 4.1.2 Methods for comparison
The selected baselines correspond to five categories. The first category is traditional machine learning models by the scikit-learn library (Pedregosa et al., 2011) using edge features, including Multilayer Perceptron (MLP) , Support Vector Machine Classifier (SVM), Logistic Regression, and Random Forest. The second category consists of GNN-based models that learn lower-order topological features from a two-way network. These models include GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024). The third category consists of Transformer-based models that utilize the Transformer architecture to perceive the entire brain picture, including Graph Transformer (Ying et al., 2021), Brain Network Transformer (Kan et al., 2022b), Tsen (Hu et al., 2023), Long-range Brain Transformer (Yu et al., 2024). The fourth category is higher-order methods based on hypergraph, including HGCN (Hao et al., 2023), HGAT (Wang et al., 2022). The fifth category is persistent homology (PH)-based models, including Brain-HORS (Santoro et al., 2024), PH-MCI (Bhattacharya et al., 2025), ATPGCN (Bian et al., 2024).
@@ 4.1.3 Implementation Details
    For GNN-based models and Transformer-based models, the correlation matrix is calculated using Pearson correlation for all time series. On this basis, the node features in the graph are the columns of the corresponding Pearson correlation matrix. The adjacency matrix in the graph is obtained by thresholding (0.5) the correlation matrix. Generally, two graph convolutional layers and an average pooling layer are used for classification.
@@
    For hypergraph-based higher-order methods, specifically, for HGCN, a graph is first constructed through the correlation matrix, where the correlation of each edge of the graph is greater than 0.7. Each node constructs 2-hyperedges, 3-hyperedges, and 4-hyperedges by selecting its k-hop neighbors in the graph. For HGAT, the feature of the central node corresponding to each hyperedge is used as the hyperedge feature. When aggregating information, the attention coefficient between the hyperedge and the node is calculated for classification. Note that more HGNN-based methods applying to brain data have been proposed recently. However, the code of these models is notpublicly available, and they require significant engineering which is hard to reproduce fairly based solely on the papers.
@@
    For the persistent homology (PH)-based models, including PH-MCI, ATPGCN. We use the persistent topological features of connected components $H_{0}$ and loops $H_{1}$) extracting by Ripserer tool. For Brain-HORS, following the method, we extract the matrix of violating triangles, the matrix of one-dimensional loops and the matrix of lower-order edge features, then flatten them and feed them into the SVM for classification.
@@
HGNN-based models HGCN 58.9$\pm$7.6 61.2$\pm$7.0 60.1$\pm$7.7 59.5$\pm$7.3
@@
HGNN-based models HGCN 55.0$\pm$20.3 58.7$\pm$23.1 55.0$\pm$20.3 52.4$\pm$20.2
@@
HGNN-based models HGCN 64.2$\pm$10.0 65.2$\pm$10.3 64.3$\pm$10.3 63.4$\pm$10.5
@@
HGNN-based models HGCN 61.0$\pm$3.0 55.2$\pm$5.3 62.3$\pm$4.6 60.4$\pm$4.3
@@ 4.2 Model Comparison
Table,2, 3, 4, 5 demonstrate the comparison results on ADNI, TaoWu, PPMI, and ABIDE datasets. The mean and standard deviation of ten-fold cross-validation were listed, with the best results highlighted in bold and the second-best results in underlining. Compared with traditional ML methods, general-purposed GNN-based models for brain network do not show a significant advantage and attain similar performance in most cases. This finding is consistent with the results reported in existing papers. This is because, under the same input conditions, neural network models typically have a larger number of parameters than conventional machine learning methods when utilizing edge features for better interpretability, making them more susceptible to overfitting. Transformer-based models predominantly outperform GNN-based models across all four datasets in most scenarios, highlighting their superior capability to capture the global structure of brain connectomes. HGNN-based models generally do not demonstrate superiority over other models, with notable performance only observed on the PPMI dataset. This phenomenon can be attributed to two main reasons: first, such models introduce excessive parameters, posing a risk of overfitting; second, their k-hop neighbor-based modeling approach fails to effectively capture the co-fluctuation characteristics among multiple brain regions, thereby inadequately representing higher-order interactions. In PH-based models, such as ATPGCN and Brain-HORs, which introduce higher-order loops or connected component features extracted via persistent homology for brain disease diagnosis, demonstrate significant superiority compared to both graph neural network (GNN)-based models and traditional machine learning models that rely solely on edge features. This highlights the necessity of incorporating higher-order topological features in brain disease diagnosis methods and provides a theoretical foundation for exploring brain diseases from higher-dimensional and more comprehensive perspectives.
@@
Our HOI-brain outperforms all 20 baselines on all datasets. In particular, it can be found that the proposed HOI-brain outperforms the traditional machine learning models on all datasets, with average accuracy improved by 12.5 (%), average precision improved by 12.9(%), average recall improved by 14.3(%) and F1-score improved by 12.7(%), respectively. And the proposed HOI-brain outperforms the GNN-based models on all datasets, with average accuracy improved by 19.6(%), average precision improved by 22.9(%), average recall improved by 19.5(%), and F1-score improved by 19.5(%), respectively. This indicates that our model, by incorporating higher-order topological features, is more beneficial for brain disease diagnosis compared to using only low-order edge features. Furthermore, the proposed HOI-Brain outperforms Transformer-based models on all datasets, with average accuracy improved by 13.0(%), average precision improved by 13.4(%), average recall improved by 10.3(%), and F1-score improved by 10.9(%), respectively. This highlights that our model, through the multi-channel brain network Transformer, can effectively and holistically integrate complementary information from heterogeneous higher-order and low-order topological features while preserving channel-specific properties. Next, when compared to HGNN-based models, HOI-Brain achieves remarkable improvements with average accuracy improved by 21.8(%), precision improved by 27.0(%), recall improved by 19.4(%), and F1-score improved by 20.5(%). These results clearly demonstrate that our proposed HOI-Brain framework significantly advances the modeling of higher-order interactions through the Multiplication of Temporal Derivatives and the extraction of topological features using persistent homology in brain networks. When specifically compared to Persistent Homology (PH) based models, HOI-Brain achieves significant improvements with average accuracy increased by 17.4(%), precision enhanced by 21.4(%), recall boosted by 14.5(%), and F1-score elevated by 16.2(%). This is partly attributed to our method linking signed HOIs in the brain to higher-order organizations, and partly due to our exploration of more complex signed quadruplet structures and voids (2D holes) formed by triangular faces and quadruplet.
@@ 4.3.1 Ablation study
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of different components on overall performance. Our first aim was to evaluate the effectiveness of more complex signed quadriplet structures and voids (2D holes) formed by triangular faces. Specifically, following the method in (Santoro et al., 2024), we extract lower-order edge features, violating triangles, one-dimensional loops. In addition, we extract unsigned good quadruplets and unsigned two-dimensional voids based our methods. We implement it within different combinations of these features on all datasets. The experimental results presented in Table 6, 7, 8, 9, indicate that integrating both higher-order and lower-order interactions proves to be more effective for brain disorder diagnosis compared to using lower-order interactions alone. Furthermore, quadruplet interactions captured by HOI-Brain have been demonstrated to be more effective in improving diagnosis than triplet interactions. This may be due to the fact that a significant proportion of triplet interactions can be decomposed into a linear combination of pairwise interactions, which is insufficient to capture the true higher-order interactions. Additionally, distinguishing between positively and negatively synergistic higher-order interactions to generate signed higher-order structures, as opposed to unsigned ones, can further improve the effectiveness of brain disease diagnosis, which may due to the fact that it can offer detailed insights into the communication within the brain.
@@
Our second aim was to evaluate the effectiveness of different components in multi-channel brain network Transformer. The different variants are: not adding signed higher-order features decoupling mechanism (wo-signed), not adding orthonormal clustering readout of multiple channels, and not adding attention-guided feature fusion mechanism (wo-fusion), and not adding orthonormal clustering readout of multiple channels(wo-cluster). From Fig.4, we can see that adding signed higher-order features decoupling mechanism and attention-guided feature fusion mechanism are always helpful as HOI-brain wins all the cases over the variant (wo-signed) or (wo-fusion). This result is consistent with our assumption that, by introducing the attention mechanism, our model can adaptively fuse positive and negative information as well as information from different channels, thereby emphasizing discriminative features while suppressing redundant information. Incorporating orthonormal clustering readout across multiple channels (wo-cluster) is highly beneficial, as this approach effectively captures both low-level and high-level feature patterns, which often reside within distinct functional modular structures.